OpenStudy (anonymous):

If 1/4 of an even positive number and 5/6 of the next larger even number have a sum of 32, what is the average of the two numbers?

OpenStudy (anonymous):

1/4(n) + 5/6(n+2) = 32 3n + 10(n+2) = 384 3n + 10n + 20 = 384 13n + 20 = 384 13n = 364 n = 28 n + 2 = 30 (28 + 30)/2 = 29

OpenStudy (anonymous):

Ah I see thank you!